integrate (x 3 +x+1)/(x 4 +x 2 +1) w.r.t x

integrate (x3+x+1)/(x4+x2+1) w.r.t x

Grade:12th Pass

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Suchita,

Split this to two integrals,

x3+x / x4 + x2 + 1 and 1/(x4+x2+1). The second integral is a standard integral (integraded by standard method).

For the first integral, you have x(x2+1)/((x^2)2+x2+1}, make the substitution x2+1 = t, and xdx = dt/2

You will have t/{(t-1)2 + t) dt = linear/Quad [again a standard integral, which one can integrate].

 

Hope that helps.

 

All the best,

Regards,

Ashwin (IIT Madras).